# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

### v1

@@ Abstract
GP models have received increasingly attentions in recent years due to their superb prediction accuracy and modeling flexibility. To address the computational burdens of GP models for large-scale datasets, distributed learning for GPs are often adopted. Current aggregation models for distributed GPs are not time-efficient when incorporating correlations between GP experts. In this work, we propose a novel approach for aggregated prediction in distributed GPs. The technique is suitable for both the exact and sparse variational GPs. The proposed method incorporates correlations among experts, leading to better prediction accuracy with manageable computational requirements. As demonstrated by empirical studies, the proposed approach results in more stable predictions in less time than state-of-the-art consistent aggregation models.
@@ 1 Introduction
Gaussian processes (GPs) (Rasmussen,, 2003) are a powerful tool for modeling and inference in various areas of machine learning, such as regression (O’Hagan,, 1978; Bishop et al.,, 1995; MacKay et al.,, 2003), classification (Kuss et al.,, 2005; Nickisch and Rasmussen,, 2008; Hensman et al.,, 2015), forecasting (Girard et al.,, 2002; Roberts et al.,, 2013), signal processing (Liutkus et al.,, 2011; Sarkka et al.,, 2013), Bayesian optimization (Snoek et al.,, 2012; Frazier,, 2018), and robotics and control (Deisenroth et al.,, 2013; Mukadam et al.,, 2016). Despite their strengths, GPs face significant challenges when applied to large-scale datasets due to the computational burden of inverting large covariance matrices.
@@
To address these challenges, numerous efficient algorithms have been proposed. Sparse approximations using $m$ inducing points (Quinonero-Candela and Rasmussen,, 2005; Titsias,, 2009; Hensman et al.,, 2013) can reduce the computational complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(m^{2}n)$ for a dataset of size $n$. Wilson and Nickisch, (2015) employed structured kernel interpolation combined with Kronecker and Toeplitz algebra to handle a large number of inducing points. Hartikainen and Särkkä, (2010); Grigorievskiy et al., (2017); Nickisch et al., (2018) reformulated GP regression as linear-Gaussian state space models, solving them with classical Kalman filtering theory. Katzfuss and Guinness, (2021); Katzfuss et al., (2022) presented a general Vecchia framework for GP predictions linear computational complexity in the total number of datasets. Chen et al., (2022); Chen and Tuo, (2022); Ding et al., (2024) represented the kernels via a sparse linear transformation to help reduce complexity. However, those approximation methods require at least $\mathcal{O}(n)$ time, which makes them impractical for very large $n$.
@@
Another direction to address the computational issue is distributed learning, which involves distributing computations across multiple units, often referred to as nodes or experts. In distributed learning (Dean et al.,, 2012), the dataset is partitioned and processed in parallel, with results aggregated to form the final model. By applying the previously mentioned GP approximation methods in distributed learning setup across $M$ units, the time complexity can be reduced to $\mathcal{O}(n/M)$. Recent aggregation models for distributed GPs (Tresp,, 2000; Hinton,, 2002; Cao and Fleet,, 2014; Deisenroth and Ng,, 2015) rely on independence assumptions and cannot offer consistent predictions, meaning the aggregated predictive distribution does not converge to the true underlying predictive distribution as the training size $n$ increases to infinity. To overcome this inconsistency, Rullière et al., (2018) introduced the nested point-wise aggregation of experts (NPAE), which incorporates covariances between the experts’ predictions but is very time-consuming. Liu et al., (2018) proposed generalized robust Bayesian committee machine (grBCM), an efficient and consistent algorithm for distributed GPs.
@@
In this work, we introduce a novel aggregation model for distributed GP to improve time complexity. Our methodology is based on GP regressions using the optimized combination technique (OptiCom) introduced in Section 3. The proposed method considers the correlations between the experts to ensure the consistency. It is also less time-consuming than NPAE and grBCM when the number of GP experts $M$ is not very large. Empirical studies show that the proposed algorithm outperforms state-of-the-art aggregation models in computational efficiency while providing stable and consistent predictions.
@@ 2 Background
This section presents the background of this paper. We begin by introducing GP regression, GP training and sparse variational GP in Section 2.1. Subsequently, we outline the methods of the training and aggregated prediction for distributed GPs in Section 2.2.
@@ 2.1.1 GP regression
A Gaussian process (GP) is a collection of random variables, any finite number of which has a multivariate normal distribution. A GP is completely defined by a mean function $\mu(\cdot)$ and a covariance (or kernel) function $k(\cdot,\cdot^{\prime})$:
$f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime})),$ (1)
where $\mu(\cdot)=\mathbb{E}[f(\cdot)]$ and $k(\cdot,\cdot^{\prime})=\mathbb{E}[(f(\cdot)-\mu(\cdot))(f(\cdot^{\prime})-\mu(\cdot^{\prime}))]=\mathrm{cov}(f(\cdot),f(\cdot^{\prime}))$.
Suppose we have a set of training points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d. noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a latent function. The GP regression imposes a GP prior over the latent function as $f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$. By the definition of GP, we would obtain a $n$-dimensional multivariate Gaussian random variable $\mathbf{f}$ by evaluating the GP $f(\cdot)$ at the set of points $\mathbf{X}=\{\mathbf{x}\}_{i=1}^{n}$,
$\mathbf{f}\sim\mathcal{N}(\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{f}}),$ (2)
where $\mathbf{f}=\big{(}f(\mathbf{x}_{1}),\cdots,f(\mathbf{x}_{n})\big{)}^{\top}$, $\bm{\mu}_{\mathbf{f}}=\big{(}\mu(\mathbf{x}_{1}),\cdots,\mu(\mathbf{x}_{n})\big{)}^{\top}$ and $\mathbf{K}_{\mathbf{f}\mathbf{f}}=\big{[}k(\mathbf{x}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{n}$ are the mean and covariance matrix obtained by evaluating at the location points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$.
In this work, we assume $f$ is a zero-mean GP, i.e., $\mu(\cdot)=0$, then the posterior distribution takes the form
$p(f|\mathbf{y})=\mathcal{N}\big{(}\mu_{f|\mathbf{y}},k_{f|\mathbf{y}}\big{)},$ (3)
with
@@
where $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}=\mathbf{K}_{\mathbf{f}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n}$ is the covariance matrix of all data points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ with diagonal observation noise, $\sigma_{\epsilon}^{2}>0$ is the noise variance, $\mathbf{I}_{n}$ is a $n\times n$ identity matrix, $\mathbf{K}_{(\cdot)\mathbf{f}}=k(\cdot,\mathbf{X})=\big{[}k(\cdot,\mathbf{x}_{i})\big{]}_{i=1}^{i=n}=k(\mathbf{X},\cdot)^{\top}=\mathbf{K}_{\mathbf{f}(\cdot)}^{\top}$ is the cross-covariance matrix.
@@ 2.1.2 GP training
The hyperparameters $\bm{\theta}$ of GPs, which may include GP variance, kernel lengthscale, are commonly learned by maximizing the log-marginal likelihood (Jones et al.,, 1998) given by
$\mathcal{L}(\bm{\theta})$ $=\log p(\mathbf{y}|\mathbf{f};\bm{\theta})=\log\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ (5)
$=-\frac{1}{2}\Big{(}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}+\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}})+n\log(2\pi)\Big{)}.$
A standard way of obtaining the optimized hyperparameters is by taking the derivative of each $\theta\in\bm{\theta}$ as follows:
$\frac{\partial}{\partial\theta}\mathcal{L}(\bm{\theta})=\frac{1}{2}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}-\frac{1}{2}\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}.$ (6)
@@ 2.1.3 Sparse variational GP (SVGP)
Exact GP regression and training requires $\mathcal{O}(n^{3})$ time complexity due to the computation of the terms $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}$, $\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1})$, and $\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}$ in eqs. 4, 5 and 6, which is prohibitive when $n$ is large. To solve this problem, Titsias, (2009) introduced sparse variational Gaussian process (SVGP), a variational framework for sparse GPs that approximates $n$ actual observations with $m$ inducing variables ($m\ll n$) by maximizing a lower bound of the true log marginal likelihood, which can reduce the time complexity of GP regression and training to $\mathcal{O}(nm^{2})$.
SVGP consists of $m$ inducing variables $\mathbf{u}=(u_{1},\ldots,u_{m})^{\top}\in\mathbb{R}^{m}$. These latent variables are values of the GP, corresponding to $m$ inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$, $\mathbf{z}_{i}\in\mathbb{R}^{d}$. We specify a Gaussian distribution over the inducing variables $q(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{S}_{\mathbf{u}\mathbf{u}})$, then integrating out $\mathbf{u}$ with $q(\mathbf{u})$ yields
[…]
where
$\mu_{f}^{\mathbf{u}}(\cdot)=\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{m}_{\mathbf{u}},$ (8a)
$k_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}(\mathbf{K}_{\mathbf{u}\mathbf{u}}-\mathbf{S}_{\mathbf{u}\mathbf{u}})\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (8b)
Here $\mathbf{K}_{(\cdot)\mathbf{u}}=\big{[}k(\cdot,\mathbf{z}_{i})\big{]}_{i=1}^{m}=\mathbf{K}_{\mathbf{u}(\cdot)}^{\top}$, and $\mathbf{K}_{\mathbf{u}\mathbf{u}}=\big{[}k(\mathbf{z}_{i},\mathbf{z}_{j})\big{]}_{i,j=1}^{m}$. The goal is to approximate the exact posterior $p(\mathbf{f},\mathbf{u}|\mathbf{y})$ by an variational distribution $q(\mathbf{f},\mathbf{u})$, so we minimize the Kullback-Leibler (KL) divergence $\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}$, which is equivalent to maximize the evidence lower bound (ELBO) defined as follows:
$\mathcal{L}_{f}:=\mathbb{E}_{q(\mathbf{f},\mathbf{u})}\left[\log\frac{p(\mathbf{y},\mathbf{f},\mathbf{u})}{q(\mathbf{f},\mathbf{u})}\right]=\log p(\mathbf{y})-\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}.$ (9)
Under a Gaussian likelihood of the form $p(\mathbf{y}|\mathbf{f})=\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$, the optimal variational distribution $\hat{q}(\mathbf{u})=\mathcal{N}(\hat{\mathbf{m}}_{\mathbf{u}},\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}})$ can be obtained by
$\hat{\mathbf{m}}_{\mathbf{u}}=\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{y},$ (10a)
$\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}}=\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{u}},$ (10b)
where $\mathbf{K}_{\mathbf{u}\mathbf{f}}=\big{[}k(\mathbf{z}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{m,n}=\mathbf{K}_{\mathbf{f}\mathbf{u}}^{\top}$, and $\mathbf{M}_{\mathbf{u}\mathbf{u}}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}}\big{]}$. The corresponding optimal ELBO then becomes
^$\hat{\mathcal{L}}_{f}=\log\mathcal{N}(\mathbf{y}|\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n})-\frac{1}{2}\sigma_{\epsilon}^{-2}\text{tr}(\mathbf{K}_{\mathbf{f}\mathbf{f}}-\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}).$ (11)
Therefore, eq. 10 can lead to the predictive distribution $\hat{q}(f)=\mathcal{N}(\hat{\mu}_{f}^{\mathbf{u}},\hat{k}_{f}^{\mathbf{u}})$ with
@@
$\hat{k}_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}(\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}-\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1})\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (12b)

### v2

@@ Abstract
GP models have received increasing attention in recent years due to their superb prediction accuracy and modeling flexibility. To address the computational burdens of GP models for large-scale datasets, distributed learning for GPs are often adopted. Current aggregation models for distributed GPs is not time-efficient when incorporating correlations between GP experts. In this work, we propose a novel approach for aggregated prediction in distributed GPs. The technique is suitable for both the exact and sparse variational GPs. The proposed method incorporates correlations among experts, leading to better prediction accuracy with manageable computational requirements. As demonstrated by empirical studies, the proposed approach results in more stable predictions in less time than state-of-the-art consistent aggregation models.
Keywords: Distributed Gaussian processes; Optimized combination technique; Aggregation models; Inducing points.
@@ 1 Introduction
Gaussian processes (GPs) (Rasmussen, 2003) are a powerful tool for modeling and inference in various areas of machine learning, such as regression (O’Hagan, 1978; Bishop, 1995), classification (Kuss et al., 2005), signal processing (Liutkus et al., 2011), and robotics and control (Deisenroth et al., 2013). Despite their strengths, GPs face significant challenges when applied to large-scale datasets due to the computational burden of inverting large covariance matrices.
@@
To address these challenges, numerous efficient algorithms have been proposed. Sparse approximations using $m$ inducing points (Titsias, 2009) can reduce the computational complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(m^{2}n)$ for a dataset of size $n$. Wilson and Nickisch (2015) employed structured kernel interpolation combined with Kronecker and Toeplitz algebra to handle a large number of inducing points. Hartikainen and Särkkä (2010) reformulated GP regression as linear-Gaussian state space models, solving them with classical Kalman filtering theory. Katzfuss and Guinness (2021) presented a general Vecchia framework for GP predictions linear computational complexity in the total number of datasets. Chen et al. (2022) represented the kernels via a sparse linear transformation to help reduce complexity. More recently, Lin et al. (2023, 2024) showed that a stochastic dual descent algorithm can efficiently yield competitive GP predictions. However, those approximation methods require at least $\mathcal{O}(n)$ time, which makes them impractical for very large $n$.
@@
Another direction to address the computational issue is distributed learning, which involves distributing computations across multiple units, often referred to as nodes or experts. In distributed learning (Dean et al., 2012), the dataset is partitioned and processed in parallel, with results aggregated to form the final model. By applying the previously mentioned GP approximation methods in distributed learning setup across $M$ units, the time complexity can be reduced to $\mathcal{O}(n/M)$. Recent aggregation models for distributed GPs (Deisenroth and Ng, 2015) rely on independence assumptions and cannot offer consistent predictions, meaning the aggregated predictive distribution does not converge to the true underlying predictive distribution as the training size $n$ increases to infinity. To overcome this inconsistency, Rullière et al. (2018) introduced the nested point-wise aggregation of experts (NPAE), which incorporates covariances between the experts’ predictions but is very time-consuming. Liu et al. (2018) proposed generalized robust Bayesian committee machine (grBCM), an efficient and consistent algorithm for distributed GPs.
@@
In this work, we introduce a novel aggregation model for distributed GP to improve time complexity. Our methodology is based on GP regressions using the optimized combination technique (OptiCom) introduced in Section 3. The proposed method considers the correlations between the experts to improve the consistency of the aggregated predictions. It is also less time-consuming than NPAE and grBCM when the number of GP experts $M$ is not very large. Empirical studies show that the proposed algorithm outperforms state-of-the-art aggregation models in computational efficiency while providing stable and consistent predictions.
@@ 2 Background
In this section, we provide the necessary background for understanding the proposed methodology. We begin with a review of Gaussian processes in Section 2.1, covering standard GP regression in Section 2.1.1, training procedures in Section 2.1.2, and sparse variational GP approximations in Section 2.1.3. We then introduce distributed GPs in Section 2.2, covering both training strategies in Section 2.2.1 and existing aggregation methods in Section 2.2.2.
@@ 2.1.1 GP regression
A Gaussian process (GP) is a collection of random variables, any finite number of which has a multivariate normal distribution. It is defined by a mean function $\mu(\cdot)$ and a kernel function $k(\cdot,\cdot)$: $f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot))$.
Suppose we have a set of training points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d. noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a latent function. The GP regression imposes a GP prior over the latent function, and we have $\mathbf{f}=f(\mathbf{X})\sim\mathcal{N}(\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{f}})$ where $\bm{\mu}_{\mathbf{f}}=\big(\mu(\mathbf{x}_{1}),\cdots,\mu(\mathbf{x}_{n})\big)^{\top}$ and $\mathbf{K}_{\mathbf{f}\mathbf{f}}=\big[k(\mathbf{x}_{i},\mathbf{x}_{j})\big]_{i,j=1}^{n}$. In this work, we assume $f$ is a zero-mean GP, i.e., $\mu(\cdot)=0$, the posterior distribution $p(f|\mathbf{y})=\mathcal{N}\big(\mu_{f|\mathbf{y}},k_{f|\mathbf{y}}\big)$ takes the form
@@
where $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}=\mathbf{K}_{\mathbf{f}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n}$ is the covariance matrix of all training points $\mathbf{X}$ with diagonal observation noise, $\sigma_{\epsilon}^{2}>0$ is the noise variance, $\mathbf{I}_{n}$ is a $n\times n$ identity matrix, $\mathbf{K}_{(\cdot)\mathbf{f}}=k(\cdot,\mathbf{X})=\big[k(\cdot,\mathbf{x}_{i})\big]_{i=1}^{i=n}=k(\mathbf{X},\cdot)^{\top}=\mathbf{K}_{\mathbf{f}(\cdot)}^{\top}$ is the cross-covariance matrix.
@@ 2.1.2 GP training
The hyperparameters $\bm{\theta}$ of GPs, which may include GP variance, kernel lengthscale, are commonly learned by maximizing the log-marginal likelihood $\mathcal{L}(\bm{\theta})=\log p(\mathbf{y}|\mathbf{f};\bm{\theta})$. A standard way of obtaining the optimized hyperparameters is by taking the derivative of each $\theta\in\bm{\theta}$.
@@ 2.1.3 Sparse variational GP (SVGP)
Exact GP regression and training as described in Section 2.1.1 and Section 2.1.2 require $\mathcal{O}(n^{3})$ time due to the computation of the terms $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}$, $\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1})$, and $\text{tr}\Big(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big)$, which is prohibitive when $n$ is large. To solve this problem, Titsias (2009) introduced sparse variational Gaussian process (SVGP), a variational framework for sparse GPs using $m\ll n$ inducing variables $\mathbf{u}=(u_{1},\ldots,u_{m})^{\top}$ with $m$ associated inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$, $\mathbf{z}_{i}\in\mathbb{R}^{d}$, which reduces the time complexity to $\mathcal{O}(nm^{2})$. The variational distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{S}_{\mathbf{u}\mathbf{u}})$ yields the predictive distribution $\hat{q}(f)=\mathcal{N}(\hat{\mu}_{f}^{\mathbf{u}},\hat{k}_{f}^{\mathbf{u}})$ as follows:
@@
$\hat{k}_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}(\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}-\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1})\mathbf{K}_{\mathbf{u}(\cdot^{\prime})},$ (2b)
where $\mathbf{K}_{(\cdot)\mathbf{u}}=\big[k(\cdot,\mathbf{z}_{i})\big]_{i=1}^{m}=\mathbf{K}_{\mathbf{u}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}\mathbf{f}}=\big[k(\mathbf{z}_{i},\mathbf{x}_{j})\big]_{i,j=1}^{m,n}=\mathbf{K}_{\mathbf{f}\mathbf{u}}^{\top}$, and $\mathbf{M}_{\mathbf{u}\mathbf{u}}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}}\big]$.
The SVGP framework introduced here is essential for our distributed learning methodology in Section 4. By employing SVGP approximations at the local expert level before aggregation, we achieve a double reduction in computational cost: first, inducing points approximation reduces the cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(nm^{2})$; second, data parallelization across $M$ experts further reduces the time complexity to $\mathcal{O}(nm^{2}/M)$.